Prove that


lim 1/(x+5)^2 = ∞


x→-5

1
Expert's answer
2022-02-27T12:45:03-0500

We should prove that "\forall M > 0 \;\; \exists \delta_M > 0 \;\; \forall x, |x+5| < \delta_M : |f(x)| > M."

So for M we should determine "\delta_M."

Let "\dfrac{1}{(x+5)^2} > M \; \Rightarrow \; (x+5)^2 < \dfrac{1}{M}, \\
|x+5| < \dfrac{1}{\sqrt{M}}\;, \delta_M = \dfrac{1}{\sqrt{M}}."

Therefore, when x tends to -5, the value of function increases and tends to infinity.